A steel rod has a length of 2 m. The temperature coefficient of linear expansion for steel is 11 X 10raised to peer 60C-1. By how much would its length change when its temperature changes by 50 °C.
1
Expert's answer
2020-12-03T10:58:41-0500

We can find the change in length of the steel rod from the formula:


"\Delta L=\alpha L\Delta T,""\Delta L=11\cdot10^{-6}\ ^{\circ}C^{-1} \cdot 2\ m\cdot 50\ ^{\circ}C=1.1\cdot10^{-3}\ m."

Answer:

"\Delta L=1.1\cdot10^{-3}\ m."